OpenStudy (anonymous):

Write an equation and solve. If four times Catherine’s age is subtracted from 5 times Jose’s age, the difference is 32 years. Jose is sixteen. Find Catherine’s age.

OpenStudy (anonymous):

Let Catherine's age be c and let Jose's age be j. Then, the world problem can be re-stated as follows: "4c - 5j = 32 and j = 16" (in years). Therefore, in years, Catherine's age is: \[ c = \dfrac {32 + 5j}{4} = \dfrac {32 + 80}{4} = \dfrac {112}{4} = \boxed {28}. \]

OpenStudy (anonymous):

Catherine's age should be 12. 5J - 4C = 32 Where J equals Jose, C equals Catherine Therefore, 80-4C = 32 C= 12.

OpenStudy (anonymous):

I stand corrected. Thanks, @Freak7820 - "subtracted from" is not what I said. Sorry and thanks again!